# When Experimental Economics Meets

### v1

@@ Abstract
Advancements in large language models (LLMs) have sparked a growing interest in measuring and understanding their behavior through experimental economics. However, there is still a lack of established guidelines for designing economic experiments for LLMs. By combining principles from experimental economics with insights from LLM research in artificial intelligence, we outline and discuss eight practical tactics for conducting experiments with LLMs. We further perform two sets of experiments to demonstrate the significance of these tactics. Our study enhances the design, replicability, and generalizability of LLM experiments, and broadens the scope of experimental economics in the digital age.
@@ 1 Introduction
The rise of large language models (LLMs), particularly since the launch of ChatGPT in November 2022, has sparked significant interest in understanding their capacities [26, 79]. Research has shown that LLMs exhibit impressive abilities beyond mere language generation, including generating computer code, solving complex math problems, and demonstrating human-like reasoning. This growing prominence has led to a surge in studies that examine how LLMs perform in economic experiments using decision-making tasks—referred to as “LLM experiments”. These studies measure preferences, rationality, heuristics, and biases of LLMs, and assess whether they can pass a form of the Turing test that evaluates behavioral similarity to humans [23, 50, 58, 66]. As LLMs continue to be integrated into various aspects of daily life and reshape our interactions with technology, these studies are becoming increasingly crucial and popular. However, despite this rapid progress, little has been done to examine the underlying rules of conducting economic experiments in the context of LLMs.
@@
In this paper, we propose eight tactics for conducting LLM experiments and perform two case studies to test the significance of these tactics. In Case Study 1, we examine the economic rationality of LLMs [23], in which LLMs make a series of budgetary decisions [4, 24, 23]. In Case Study 2, we examine the economic preferences of LLMs [66], in which LLMs make decisions in a series of behavioral games, including the dictator game, ultimatum game, public goods game and bomb risk game. We conduct the experiments using four popular LLMs, including GPT-4o-2024-11-20 (hereafter referred to as GPT), DeepSeek-V3-0324 (DeepSeek), Llama-3.1-8B-Instruct (Llama) and Qwen2.5-7B-Instruct (Qwen). We find that assigned personas, such as occupation, significantly affect preferences but not rationality. Second, the single-turn dialogue approach significantly decreases the rationality of Llama and Qwen, while that of GPT and DeepSeek is not affected by the dialogue type. Third, limiting responses to multiple-choice options instead of allowing open-ended responses reduces the rationality of Llama and Qwen. It also significantly changes the outputs of all four LLMs in approximately half of the behavioral games. In general, the results of both case studies shed light on the significance of design, replicability, and generalizability in LLM experiments and support the validity of our proposed tactics.
@@ 2 Experimental Economics: from Humans to LLMs
In this section, we first review the procedures of existing research on LLM experiments. Then, combining insights from both experimental economics and LLM studies in artificial intelligence, we formally propose eight tactics for conducting LLM experiments.
In addition, we provide some background knowledge in the appendices. In particular, we summarize several key principles for conducting human experiments by reviewing classical textbooks and survey papers [30, 41, 28]. In the experimental design stage, these include (1) follow the induced value theory, (2) avoid deception, and (3) minimize experimenter demand effects. In the experimental implementation stage, the principles are (1) use context-neutral instructions, (2) make instructions comprehensible, (3) provide illustrative examples, (4) incorporate comprehension assessments, and (5) prevent inter-subject communication. We detail these principles in SI Appendix, Section A. Moreover, we also provide a comprehensive overview of the development of LLMs, introduce various types of LLMs, and outline the procedures to prompt them in SI Appendix, Section B.
2.1 LLM Experiments: An Emerging Era
We summarize 12 papers involving LLM experiments that have been published in leading journals or conferences (SI Appendix, Table S1), including Proceedings of the National Academy of Sciences (PNAS), Nature Human Behaviour, Marketing Science, and Manufacturing & Service Operations Management, among others. Although there is a rapidly growing literature on this topic [e.g., 58, 94, 31], we decide to include only published papers because their experimental procedures are finalized and many of them provide the code which enables us to replicate their results. We apply what we have learned from both artificial intelligence and human experiments to evaluate these studies. We find that these studies exhibit a large degree of heterogeneity in the choice of parameters in the prompts, as summarized below (SI Appendix, Table S2).
@@ 2.1 LLM Experiments: An Emerging Era
    Temperature. Temperature is a unique design parameter raised by LLM experiments. Take the models in GPT family as an example, three out of twelve studies set temperature as 0, five choose 1, one chooses 0.7, and two studies have tried multiple values such as 0 and 1 and do not find significant differences across temperatures [23, 55].
@@
    Persona (who) matters. Prior experimental studies have shown that individual choices are affected by their demographics [29, 33]. Although most LLM experiments evaluate the performance of LLMs without assigning specific demographics, some examine the impact of demographics with different purposes. To compare with the choices of undergraduate students in the games, Brookins and DeBacker [13] explicitly assign LLMs the persona of an undergraduate student. To explore whether LLMs can exhibit human-like behavioral heterogeneity across groups, Chen et al. [23] and Mei et al. [66] vary different demographics such as gender, ethnicity, and occupation.
@@
    Illustrative example and understanding question. Nearly all LLM experiments do not include examples and understanding questions. We conjecture this is because the tasks used in these studies are simple and easy to understand. However, we notice the use of illustrative examples [66, 88] and understanding questions [23]. Take the latter as an example, the rationale lies in its goal to compare the performance of LLMs with human subjects, where understanding questions are employed in human experiments.
@@
    Single-turn vs. multi-turn dialogue. Among these studies, four of them involve situations that each LLM agent needs to answer multiple questions [7, 23, 89, 66], while the remaining studies ask each agent only to answer one question. In practice, both single-turn dialogue and multi-turn dialogue can be applied to instruct LLMs to collect answers for multiple questions. SI Appendix, Figure S1 provides an illustration of these methods. Within these three studies, one utilizes the single-turn dialogue, while the three two employ the multi-turn dialogue. Notably, Mei et al. [66] pose multiple questions to each LLM agent and provide feedback to elicit its strategy in several games. Such feedback can only be delivered using the multi-turn dialogue, thus single-turn dialogue is not feasible due to the experimental design.
@@
    Invalid answer. There are instances in which the models fail to generate valid answers. For example, LLMs may produce statements such as “As an AI language model, I cannot participate in surveys nor accept tokens as payment” [44] or “As an AI language model, I am not capable of making decisions on my own…” [23]. Another type of invalid answers features that LLMs generate elaborative explanatory content without explicitly providing the requested outputs. In handling these invalid responses, four studies report the proportion of them in the analysis, whereas the others do not mention the proportion.
2.2 Eight Tactics for LLM Experiments
Building on key principles from human experiments and the experimental protocols in previous LLM experiments, below we present eight key tactics for conducting LLM experiments. First, these tactics aim to align the procedures in LLM experiments with those in human experiments as closely as possible, thus facilitating the comparison in experimental results between LLM agents and human subjects. Second, with the burgeoning of LLM experiments, we hope these tips can be useful in improving the replicability of LLM experimental results. In the following, we will discuss each tactic by integrating insights from both experimental methods in economics and prompt engineering in artificial intelligence.
2.2.1 Tactics in LLM Experimental Design
Temperature. The selection of temperature involves the trade-off between adopting the solution with higher probability and exploring a more broad solution space [78]. Human participants often exhibit a certain degree of heterogeneity. To mimic their behavior, we should allow the variation of responses between LLM agents. Meanwhile, studies have shown that there are no significant behavioral differences when different temperatures are used [23, 55, 73, 78]. Therefore, we suggest using the default temperatures recommended by model providers. When there are no default values, setting the temperature at 1 is a natural choice. This value does not scale the logits (see SI Appendix, Section B.3), thus best mirroring the application of softmax during the training process, which does not involve the temperature parameter [76]. In addition, some model providers allow users to specify a random seed parameter. To enhance the replicability of results, researchers can set the seed when prompting LLMs via the application programming interface (API). We summarize these discussions in the following.
Tactic 1: Use the default temperature and fix the random seed.
Persona. Following the conventions of experimental economics, where experimenters do not explicitly recruit participants with specific demographics unless this is part of the research design, we recommend not including specific personas in the prompts. Instead, LLMs should be regarded as counterparts to representative individuals from the general human population. However, if the goal is to assess whether LLMs with diverse demographic characteristics respond differently to tasks [23, 66], we can implement it by assigning different personas. The caveat here is that the comparison in results between such studies should be cautious due to the differences in selected tasks, evaluated models, and persona types.
Sample size. Due to the requirements of statistical analysis, determining the sample size constitutes an important step in LLM experiments. To assess the performance of LLMs, each instance of prompting an LLM with a fresh memory state is usually treated as an independent observation. Conceptually, each observation serves as a realization drawn from the performance distribution. Conditional on this notion of an observation, the determination of sample size in LLM experiments should adhere to the same principles as in human experiments. The central criterion remains statistical power analysis, which determines the minimum required sample size to detect the hypothesized treatment effect with a pre-specified level of power ($\beta$), i.e., committing type II error with the probability of $1-\beta$, given the significance level and standard deviation of the targeted outcome.
Tactic 3: Determine sample size according to the standard procedure in experimental economics.
Incentive. Based on induced value theory [82, 83], the use of monetary incentives has become a standard practice in laboratory experiments, although the extent to which human behavior differs under real versus hypothetical incentives remains inconclusive [e.g., 48, 56, 11, 43]. Since real monetary rewards cannot be applied to LLMs, the feasible alternative is to instruct LLMs to maximize their hypothetical payoffs. Therefore, clear payment rules should be included in the prompts as in the instructions for human participants. However, unlike humans, LLMs lack an intrinsic utility function; their data generation is driven by externally imposed optimization rules in training processes. This indicates that LLMs rarely report untruthfully that incentives are designed to address. Note that although LLMs may exhibit hallucination characterized by nonsensical or unfaithful responses, this phenomenon results from the training data, training processes, and decoding strategies [51] rather than strategic deception. Based on these facts, we conjecture that LLMs’ performance may change little when such hypothetical incentives are removed.
Tactic 4: Use the typical payment scheme to be comparable to human experiments.
2.2.2 Tactics in LLM Experimental Implementation
Tactic 5: Incorporate examples and comprehension questions especially for complex tasks.
Dialogue type. Since the structure of multi-turn dialogue is more similar to the sequential decision-making process of human subjects, we suggest using it for tasks involving multiple questions for a single agent, rather than the single-turn dialogue which has advantage in the efficiency of model inputs.
Tactic 6: Use multi-turn dialogue for tasks with multiple questions.
Answer type. The construction of the answer format can influence model performance. For example, providing Likert scales or multiple choices may activate memory in the model training corpus, or enable models to infer experimental purposes and act in a demanding way [55]. Furthermore, discretizing continuous feasible sets has been shown to significantly decrease the performance of the model [23]. Consequently, the open-ended answer format is recommended rather than multiple-choice, unless the latter format is required by the nature of tasks. For example, the Linda problem requires multiple options with inclusion relationships to measure the conjunction fallacy [52].
Tactic 7: Use an open-ended answer format whenever feasible.
Invalid answer. In prior LLM experiments, the common practice is to exclude invalid answers from the dataset and continue data collection until the target sample size is reached [e.g., 23, 13, 44, 66]. However, the frequency of invalid answers also provides valuable information on model performance [64, 22]. Therefore, although the content of invalid answers may remain unused, we believe that the proportion of such responses should be documented and reported as a performance metric.
Tactic 8: Report the proportion of invalid responses.
@@ 3 Case Study 1: LLMs’ Economic Rationality
In this section, we implement the budgetary tasks of Chen et al. [23] to examine the extent to which the tactics discussed above influence the performance of LLMs.
@@ 3.1 Experimental Design
We choose four representative LLMs to carry out our tasks: (1) GPT-4o-2024-11-20, the advanced LLM developed by OpenAI, which is widely regarded as the pioneering leader in the field, (2) DeepSeek-V3-0324, the open-sourced model with the largest parameter size [65], (3) Llama-3.1-8B-Instruct, a widely studied model in artificial intelligence and the most downloaded text-generation model released since 2024, ranked by Hugging Face, and (4) Qwen2.5-7B-Instruct, an emerging model that has gained increasing attention in the field of artificial intelligence since early 2025 [e.g., 42, 68, 93]. All four models are prompted via the API using Python.
@@ 3.1.1 Budgetary Decision Tasks
Economic rationality is a central assumption in economics and is widely evaluated through various domains of decision-making [e.g., 4, 24, 25, 35, 54]. We use standard budgetary decision tasks in risk preference [24] and social preference [4] to measure economic rationality. In the risk preference domain, the agent allocates 100 points to two accounts, where the exchange rates between points and payoffs differ. She receives the reward from a random account with probability 50%. For the social preference, the agent allocates 100 points between herself and another randomly matched agent. These points are converted into payoffs with distinct exchange rates for the two agents. We construct 25 rounds of tasks in each simulation, with budget lines that vary between rounds.
@@ 3.1.2 Baseline Condition
In the baseline condition, we closely follow our summarized tactics to execute the experiment. Specifically, the temperature is set to the default value of each LLM (Tactic 1), [^1] The default temperatures are 1 for GPT (https://platform.openai.com/docs/api-reference/responses/create), 0.3 for DeepSeek (https://huggingface.co/deepseek-ai/DeepSeek-V3-0324), 0.6 for Llama (https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct/blob/main/generation_config.json) and 0.7 for Qwen (https://huggingface.co/Qwen/Qwen2.5-7B-Instruct/blob/main/generation_config.json), all last accessed on May 25, 2025. and there is no specific persona included in the system message (Tactic 2). We outline the payment rule in which one round is randomly selected for payment, to hypothetically incentivize the LLMs (Tactic 4). In the user message, we also provide a demonstration example and evaluate the degree of comprehension using understanding questions (Tactic 5). In each simulation, we use a multi-turn approach, where each round of question is asked separately, with the preceding dialogue included (Tactic 6). Thus, each LLM makes one decision per round. Additionally, we use the open-ended answer type (Tactic 7) and restrict the output to the JSON format to streamline data collection. When invalid answers occur, we exclude them and continue until 25 valid answers are collected in a simulation. [^2] Invalid answers include instances where (1) LLMs refuse to provide answers, (2) the outputs do not adhere to the specified formats, i.e., JSON format in Case Study 1 or number highlighted using “[[]]” in Case Study 2, and (3) the answers fail to meet task requirements, e.g., the total points allocated to two accounts do not sum to 100 in Case Study 1 or the number of points falls outside the specified range in both case studies. Meanwhile, the proportion of invalid answers is also documented (Tactic 8). We repeat the above procedure 100 times to generate 100 simulations, ensuring sufficient statistical power for our subsequent comparisons between conditions (Tactic 3). [^3] By assuming equal variance and specifying the minimum detectable difference in CCEI scores (see Section 3.2) of 0.01, we use the standard deviation obtained in Chen et al. [23] and compute the minimum sample size for each condition, which is 24 in the risk preference and 18 in the social preference, respectively.

### v2

@@ Abstract
Advancements in large language models (LLMs) have sparked a growing interest in measuring and understanding their behavior through experimental economics. However, there is still a lack of established guidelines for designing economic experiments for LLMs. Inspired by principles from experimental economics with insights from LLM research in artificial intelligence, we outline key considerations in the experimental design and implementation stage, and perform two sets of experiments to assess the impact of these considerations on LLMs’ responses. Based on our findings, we discuss seven practical tactics for conducting experiments with LLMs. Our study enhances the design, replicability, and generalizability of LLM experiments, and broadens the scope of experimental economics in the digital age.
@@ 1 Introduction
The rise of large language models (LLMs), particularly since the launch of ChatGPT in November 2022, has sparked significant interest in understanding their capacities [26, 82]. Research has shown that LLMs exhibit impressive abilities beyond mere language generation, including generating computer code, solving complex math problems, and demonstrating human-like reasoning. This growing prominence has led to a surge in studies that examine how LLMs perform in economic experiments using decision-making tasks—referred to as “LLM experiments”. These studies measure the preferences, rationality, heuristics, and biases of LLMs, and assess whether they can pass a form of the Turing test that evaluates behavioral similarity to humans [23, 51, 60, 68]. As LLMs continue to be integrated into various aspects of daily life and reshape our interactions with technology, these studies are becoming increasingly crucial and popular [69]. However, despite this rapid progress, little has been done to examine the underlying rules of conducting economic experiments in the context of LLMs.
@@
In this paper, we design two sets of experiments to examine the effect of some key considerations in experimental protocols on LLMs’ responses from different perspectives. In Case Study 1, we examine the economic rationality of LLMs [23, 56], in which LLMs make a series of budgetary decisions [4, 24, 23]. In Case Study 2, we examine the economic preferences of LLMs [68, 97], in which LLMs make decisions in a series of behavioral games, including the dictator game, ultimatum game, public goods game and bomb risk game. We conduct the experiments using four popular LLMs, including GPT-4o-2024-11-20 (hereafter referred to as GPT), DeepSeek-V3-0324 (DeepSeek), Llama-3.1-8B-Instruct (Llama) and Qwen2.5-7B-Instruct (Qwen). We find that assigned personas, such as occupation, significantly affect preferences but not rationality. Second, the single-turn dialogue approach significantly decreases the rationality of Llama and Qwen, while that of GPT and DeepSeek is not affected by the dialogue type. Third, limiting responses to multiple-choice options instead of allowing open-ended responses reduces the rationality of Llama and Qwen. It also significantly changes the outputs of all four LLMs in approximately half of the behavioral games. The results of two case studies inform actionable tactics, and shed light on the significance of design, replicability, and generalizability in LLM experiments. In a broad sense, our study contributes to the replicability of experimental studies by providing a framework that can be applied consistently and be comparable across experiments, thereby strengthening the reliability and validity of experimental findings [15, 16].
@@ 2 Experimental Economics: from Humans to LLMs
In this section, we begin by outlining foundational principles drawn from the practices of human experiments within experimental economics, then review the procedures of existing research on LLM experiments. In particular, we summarize several key principles for conducting human experiments by reviewing classical textbooks and survey papers [30, 42, 28]. In the experimental design stage, these include (1) follow the induced value theory, (2) avoid deception, and (3) minimize experimenter demand effects. In the experimental implementation stage, the principles are (1) use context-neutral instructions, (2) make instructions comprehensible, (3) provide illustrative examples, (4) incorporate comprehension assessments, and (5) prevent inter-subject communication. We detail these principles in Supplementary Section A. Moreover, we also provide a comprehensive overview of the development of LLMs, introduce various types of LLMs, and outline the procedures to prompt them in Supplementary Section B.
Next, we summarize 12 papers involving LLM experiments that have been published in leading journals or conferences (Supplementary Table 1), including Proceedings of the National Academy of Sciences (PNAS), Nature Human Behaviour, Marketing Science, and Manufacturing & Service Operations Management, among others. Although there is a rapidly growing literature on this topic [e.g., 60, 56, 97, 31], we decide to include only published papers because their experimental procedures are finalized and many of them provide the code which enables us to replicate their results. We apply what we have learned from both artificial intelligence and human experiments to evaluate these studies. We find that these studies exhibit a large degree of heterogeneity in the choice of parameters in the prompts (Supplementary Table 2), as summarized below.
@@
    Temperature. Temperature is a unique design parameter raised by LLM experiments. Take the models in GPT family as an example, three out of twelve studies set temperature at 0, five choose 1, one chooses 0.7, and two studies have tried multiple values such as 0 and 1 and do not find significant differences across temperatures [23, 57].
@@
    Persona. Prior experimental studies have shown that individual choices are affected by their demographics [29, 33]. Although most LLM experiments evaluate the performance of LLMs without assigning specific demographics, some examine the impact of demographics with different purposes. To compare with the choices of undergraduate students in the games, Brookins and DeBacker [13] explicitly assign LLMs the persona of an undergraduate student. To explore whether LLMs can exhibit human-like behavioral heterogeneity across groups, Chen et al. [23] and Mei et al. [68] vary different demographics such as gender, ethnicity, and occupation.
@@
    Incentive. Four studies incorporate a hypothetical incentive scheme into the prompt, while six do not. The remaining two studies apply the incentive to some tasks and omit it in others.
@@
    Illustrative example and understanding question. Nearly all LLM experiments do not include examples and understanding questions. We conjecture this is because the tasks used in these studies are simple and easy to understand. However, we notice the use of illustrative examples [68, 91] and understanding questions [23] in some studies. Take the latter as an example, the rationale lies in its goal to compare the performance of LLMs with human subjects, where understanding questions are employed in human experiments.
@@
    Single-turn vs. multi-turn dialogue. Among these studies, four of them involve situations that each LLM agent needs to answer multiple questions [7, 23, 92, 68], while the remaining studies ask each agent only to answer one question. In practice, both single-turn dialogue and multi-turn dialogue can be applied to instruct LLMs to collect answers for multiple questions. Supplementary Figure 1 provides an illustration of these methods. Within these three studies, one utilizes the single-turn dialogue, while the three two employ the multi-turn dialogue. Notably, Mei et al. [68] pose multiple questions to each LLM agent and provide feedback to elicit its strategy in several games. Such feedback can only be delivered using the multi-turn dialogue, thus single-turn dialogue is not feasible due to the experimental design.
  6. 6.
@@
    Invalid answer. There are instances in which models fail to generate valid answers. For example, LLMs may produce statements such as “As an AI language model, I cannot participate in surveys nor accept tokens as payment.” [45] or “As an AI language model, I am not capable of making decisions on my own.” [23]. Another type of invalid answers features that LLMs generate elaborative explanatory content without explicitly providing the requested outputs. In handling these invalid responses, four studies report the proportion of them in the analysis, whereas the others do not mention the proportion.
@@ 3 Case Study 1: LLMs’ Economic Rationality
In this section, we implement the budgetary tasks of Chen et al. [23] to examine the extent to which the considerations discussed above influence the performance of LLMs.
@@ 3.1 Experimental Design
We choose four representative LLMs to carry out our tasks: (1) GPT-4o-2024-11-20, the advanced LLM developed by OpenAI, which is widely regarded as the pioneering leader in the field, (2) DeepSeek-V3-0324, the open-sourced model with the largest parameter size [67], (3) Llama-3.1-8B-Instruct, a widely studied model in artificial intelligence and the most downloaded text-generation model released since 2024, ranked by Hugging Face, and (4) Qwen2.5-7B-Instruct, an emerging model that has gained increasing attention in the field of artificial intelligence since early 2025 [e.g., 43, 71, 96]. All four models are prompted via the application programming interface (API) using Python.
@@ 3.1.1 Budgetary Decision Tasks
Economic rationality is a central assumption in economics and is widely evaluated through various domains of decision-making [e.g., 4, 24, 25, 35, 55]. We use standard budgetary decision tasks in risk preference [24] and social preference [4] to measure economic rationality. In the risk preference domain, the agent allocates 100 points to two accounts, where the exchange rates between points and payoffs differ. She receives the reward from a random account with probability 50%. For the social preference, the agent allocates 100 points between herself and another randomly matched agent. These points are converted into payoffs with distinct exchange rates for the two agents. We construct 25 rounds of tasks in each simulation, with budget lines that vary between rounds.
@@ 3.1.2 Baseline Condition
In the baseline condition, we closely align the experimental procedures with those commonly used in human experiments. The temperature is set to the default value of each LLM, [^1] The default temperatures are 1 for GPT (https://platform.openai.com/docs/api-reference/responses/create), 0.3 for DeepSeek (https://huggingface.co/deepseek-ai/DeepSeek-V3-0324), 0.6 for Llama (https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct/blob/main/generation_config.json) and 0.7 for Qwen (https://huggingface.co/Qwen/Qwen2.5-7B-Instruct/blob/main/generation_config.json), all last accessed on May 25, 2025. and there is no specific persona included in the system message. We outline the payment rule in which one round is randomly selected for payment to hypothetically incentivize LLMs. In the user message, we also provide a demonstration example and evaluate the degree of comprehension using understanding questions. In each simulation, we use a multi-turn approach, where each round of question is asked separately, with the preceding dialogue included. Thus, each LLM makes one decision per round. Additionally, we use the open-ended answer type and restrict the output to the JSON format to streamline data collection. When invalid answers occur, we exclude them and continue until 25 valid answers are collected in a simulation. [^2] Invalid answers include instances where (1) LLMs refuse to provide answers, (2) the outputs do not adhere to the specified formats, i.e., JSON format in Case Study 1 or number highlighted using “[[]]” in Case Study 2, and (3) the answers fail to meet task requirements, e.g., the total points allocated to two accounts do not sum to 100 in Case Study 1 or the number of points falls outside the specified range in both case studies. Meanwhile, the proportion of invalid answers is also documented. We repeat the above procedure 100 times to generate 100 simulations, ensuring sufficient statistical power for our subsequent comparisons between conditions. [^3] By assuming equal variance and specifying the minimum detectable difference in CCEI scores of 0.01, we use the standard deviation obtained in Chen et al. [23] and compute the minimum sample size for each condition, which is 24 in the risk preference and 18 in the social preference, respectively.